This is a General Factoring Section...
Find Solutions?
y(2y+7)=15

Answer by scott8148(6628)

2y^+7y=15

subtracting 15 __ 2y^2+7y-15=0

factoring __ (2y-3)(y+5)=0

2y-3=0 __ y=3/2

y+5=0 __ y=-5